OpenStudy (anonymous):

If you save one penny on January 1, two pennies on January 2, three pennies on January 3, and continue this pattern for one year (not a leap year), what will be the value of your entire savings, in dollars, at the end of that one year? Express your answer as a decimal.

OpenStudy (atlas):

I think you know the answer...........you just don't know the number right?

OpenStudy (atlas):

The answer is Total money (in pennies) = 1+2 +3 +4+..........................+364+365 right

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

What equation would i use to figure this out?

OpenStudy (anonymous):

I think this involves the 365! equation...just keep adding :P

OpenStudy (anonymous):

That's a lot of adding though! D:

OpenStudy (agent0smith):

You can do this with Gaussian addition http://mathcentral.uregina.ca/QQ/database/QQ.02.06/jo1.html, or do it as an arithmetic sequence

OpenStudy (agent0smith):

Have you learned arithmetic sequences...?

OpenStudy (anonymous):

Yes, I've learned about them. I just didn't remember which formula to use. I will try the Gaussian addition though.

OpenStudy (anonymous):

It's 66,795 pennies, now you just have to decide how many dollars thats worth :)

OpenStudy (atlas):

sry I was away: let Total Amount be S So S = 1+2+3+................................+364+365 now reverse it S = 365 +364 +..............................+2+1 Add them up and see........you don't need any formula for this :0

OpenStudy (agent0smith):

If you use arithmetic sequences, you'll get the gaussian addition formula anyway

OpenStudy (atlas):

Add S forward and S reverse

OpenStudy (anonymous):

its 66795 pennies, move the decimal place voer 2 and you get 667.95 dollars.

OpenStudy (anonymous):

I used the Gaussian addition formula and I got 66,795 pennies. So I guess $667.95 is the answer?

OpenStudy (anonymous):

yes tegiebear :D

OpenStudy (anonymous):

Okay, thank you guys very much! :)

OpenStudy (atlas):

S = 1 + 2 + 3 + 4 +.....................................364 + 365 S =365+ 364+ 363+362+.........................................2 + 1 Adding 2S = 366 + 366 + 366 +...................365 times 2S = 365*366 S = 365*366/2 Thought this trick will help you

OpenStudy (agent0smith):

@atlas the link i first posted shows that exact method (Gaussian addition)

OpenStudy (atlas):

ah sry i didn't see it............... :P

OpenStudy (agent0smith):

no prob, plus it was only using 1 to 100 not 1 to 365

OpenStudy (atlas):

oic